OpenStudy (anonymous):

Can someone PLEASE help me complete this table of values? heres the function: f(x)=8/x^3. there is already numbers listed for the x slot, just not the other. The numbers are -4,-3,-2,-1,0,1,2,3,4. There asking me to fill out the rest and I dont know how, Someone please help me out here.

OpenStudy (anonymous):

Anybody here?

OpenStudy (anonymous):

Hi!

OpenStudy (anonymous):

If you have posted question here, then you will get its solution here only. :P

OpenStudy (anonymous):

I will help you only on one condition.. Agree?

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

you are to just interact with me, whatever I am telling you, asking you, just read my comments and answer honestly, if you know or don't know, both are same to me, just reply when you are asked to do so. Any doubt, anywhere, stop me in between and ask me there itself? Okay now?

OpenStudy (anonymous):

Of course I will ! thank you so much!

OpenStudy (anonymous):

Let us start with our function, see the other openstudy members are studying, and we are attending a function. :)

OpenStudy (anonymous):

The function which we are attending is : \[f(x) = \frac{8}{x^3}\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So, let us draw a table that your question is requiring us to: 

OpenStudy (anonymous):

Okay, and don't laugh, that is my table. :P

OpenStudy (anonymous):

Its beautiful :p

OpenStudy (anonymous):

So, you are given with some \(x\) values.. Let us take first one: \(x =-4\), right?

OpenStudy (anonymous):

Now, you are just plug in the value of \(x\) given to you and then find out \(f(x)\) corresponding to that \(x\) value.

OpenStudy (anonymous):

Yes. -4,-3,-2,-1, 0, 1, 2, 3,4

OpenStudy (anonymous):

*you are to just plug.......

OpenStudy (anonymous):

Like first one I did: \[f(x) = \frac{8}{x^3} \\ f(-4) = \frac{8}{(-4)^3}\] Getting my this step?

OpenStudy (anonymous):

Yes. So far so good.

OpenStudy (anonymous):

Now tell me : What is the value of \((-4)^3\) = ?

OpenStudy (anonymous):

-64.

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

\[f(-4) = \frac{8}{-64} \implies \color{blue}{f(-4) = \frac{-1}{8}}\]

OpenStudy (anonymous):

so -1/8 goes in the first box?

OpenStudy (anonymous):

Or in decimal form: \[\color{green}{f(-4) = -0.125}\]

OpenStudy (anonymous):

Made sense?

OpenStudy (anonymous):

yep, or you can write -0.125 as well. It is your wish.

OpenStudy (anonymous):

Will it be okay if i just put -1/8 thought?:)

OpenStudy (anonymous):